OpenStudy (anonymous):

what is the slope and the intercept of 12x-3y=15

OpenStudy (abmon98):

-3y=-12x+15 y=4x-5 y=mx+c m=+4 c=-5

OpenStudy (anonymous):

First step is to get y by itself. To do that move the 12x to the other side by subtracting it. So you have: -3y=-12x+15 Now to get rid of the -3 in front of the y you have to divide it from both sides. \[\frac{ -3y }{-3 }=\frac{ -12x }{ -3 }+\frac{ 15 }{-3 }\] The -3's cancel and y is by itself. The -12x becomes a 4x and the 15 becomes -5. So you have: y = 4x -5 The 4 is the slope and the -5 is the intercept.